OpenStudy (yrivers36):

An airliner on a coast to coast trip flying at at constant air speed covers the distance from Pittsburgh to Chicago, a distance of 600 miles, in 1 hour. On the return trip, flying at the same air speed and with the wind conditions unchanged the same distance is covered in 50 minutes. What is the air speed of the plane, what is the wind speed in the west to east direction? Please with explanation.

OpenStudy (rhystic):

set up a cross multiplication problem 600 miles / 50 minutes = x miles / 60 min to get miles per hour of the trip back , this yields 600*60/50 = 720 mph this is the new airspeed of the aircraft to travel 600 miles in 50 minutes.

OpenStudy (anonymous):

m-w=600 m+w=720

OpenStudy (anonymous):

2w=120 w=60 m=660

OpenStudy (yrivers36):

why is it 2w